Solve the equation by completing the square.
(b-5)(b-1)=6
Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.
The solution set it ____

Answer by jorel1380(3719)

(b-5)(b-1)=6
b²-6b+5=6
b²-6b+5+4=6+4=10
(b-3)²=10
b-3=+/-√10
b=3+/-√10. ☺☺☺☺